I need answer to a reasoning question:

If TREE is coded as 91 and FEED as 47, then MEET is coded as?

Please tell me solution.

Grade:Select Grade

1 Answers

ng29
8 years ago
answer is 118
there is a pattern in this question
TREE               20*1 + 18*2 + 5*3 5*4 = 20+36+15+20 = 91  
FEED               6*1 + 5*2 + 5*3 + 4*4 = 6+10+15+16 = 47
MEET             13*1 +  5*2  +  5*3  +  20*4  = 13+10+15+80 =118
i will explain u for TREE 
RANKING OF T IN ALPHABETICAL ORDER IS 20     AND  1  PLACE IN WORD
RANKING OF R IN ALPHABETICAL ORDER IS 18       AND  2  PLACE IN WORD
RANKING OF E IN ALPHABETICAL ORDER IS 5           AND  3  PLACE IN WORD
RANKING OF E IN ALPHABETICAL ORDER IS 5              AND  4  PLACE IN WORD
 
and code is given as sum of multiplication of alphabet places in alphabetical order and in given word order
 
APPROVE IF USEFUL 
APPROVE IF USEFUL 
APPROVE IF USEFUL